### v1

# Self-Organizing Dual-Buffer Adaptive Clustering Experience Replay (SODASER) for Safe Reinforcement Learning in Optimal Control

Abstract
This paper proposes a novel reinforcement learning framework, named Self-Organizing Dual-buffer Adaptive Clustering Experience Replay (SODACER), designed to achieve safe and scalable optimal control of nonlinear systems. The proposed SODACER mechanism consisting of a Fast-Buffer for rapid adaptation to recent experiences and a Slow-Buffer equipped with a self-organizing adaptive clustering mechanism to maintain diverse and non-redundant historical experiences. The adaptive clustering mechanism dynamically prunes redundant samples, optimizing memory efficiency while retaining critical environmental patterns. The approach integrates SODASER with Control Barrier Functions (CBFs) to guarantee safety by enforcing state and input constraints throughout the learning process. To enhance convergence and stability, the framework is combined with the Sophia optimizer, enabling adaptive second-order gradient updates. The proposed SODACER-Sophia’s architecture ensures reliable, effective, and robust learning in dynamic, safety-critical environments, offering a generalizable solution for applications in robotics, healthcare, and large-scale system optimization. The proposed approach is validated on a nonlinear Human Papillomavirus (HPV) transmission model with multiple control inputs and safety constraints. Comparative evaluations against random and clustering-based experience replay methods demonstrate that SODACER achieves faster convergence, improved sample efficiency, and a superior bias-variance trade-off, while maintaining safe system trajectories, validated via the Friedman test.
keywords:
Adaptive Clustering, Dual-Buffer Experience Replay, HPV model, Nonlinear Optimal Control, Safe Reinforcement Learning.
Introduction
The optimal control of nonlinear, continuous-time systems under state and input constraints remains a critical challenge in control theory, especially for high-dimensional, complex systems where safety and performance are paramount. Developing control strategies for these systems demands approaches that balance exploration and exploitation adaptively, ensuring that constraints are respected under varied operational conditions. In this context, Reinforcement Learning (RL) has emerged as a powerful framework, offering adaptive, data-driven solutions that optimize system performance while navigating uncertainties and constraints. This adaptability makes RL particularly suited for scenarios where traditional model-based approaches struggle with generalization and dynamic changes in system behavior [1, 2].
Control Barrier Functions (CBFs) are widely used to ensure safe operation within constrained control systems, guiding the system away from unsafe states by enforcing state-dependent safety limits. CBFs have been successfully integrated into RL algorithms, enabling safe exploration and optimal performance within RL frameworks. This integration has become especially valuable in applications requiring both safety and adaptability in real-time [3, 4].
While RL has shown success across various domains, achieving safe, optimal control for nonlinear systems with state and input constraints remains challenging. Techniques such as Lyapunov-based methods and constraint-satisfaction criteria have been incorporated into RL to address safety concerns [5, 6, 7]. However, these approaches face limitations when applied to high-dimensional systems with rapidly evolving dynamics. Furthermore, they often struggle with the bias-variance trade-off inherent in RL, where prioritizing stability may reduce adaptability. Experience Replay (ER) has become instrumental in addressing these RL challenges, enhancing learning efficiency and adaptability by retaining and reusing past interactions [8].
Experience replay fundamentally improves RL efficiency by storing past experiences, which reduces the need for continuous system interaction a significant benefit for applications where real-time experimentation is impractical [9, 10]. While traditional ER methods utilize random sampling, these approaches can fall short in nonstationary environments or complex control tasks. Schaul et al. [11] proposed Prioritized Experience Replay (PER), which enhances sample efficiency by assigning higher sampling probabilities to experiences with larger temporal-difference errors. Although PER improves focus on crucial transitions, it risks overemphasizing outliers or outdated samples, potentially leading to instability in changing environments.
Recent advancements in ER have focused on enhancing sample diversity and relevance through clustering mechanisms. Selective experience replay, proposed by Isele and Cosgun [12], balances recent and high-value experiences to mitigate catastrophic forgetting in continual learning. Zhang and Sutton [13] further explored deep RL with clustering-enhanced ER, showing that clustering improves sample relevance and reduces memory demands. However, clustering-based ER methods often struggle to balance the trade-off between recent and historical data, and typically lack the adaptability required for rapidly evolving environments.
The dual-buffer approach with clustering presents a novel solution to these limitations by offering both short-term adaptability and long-term experience diversity. The Fast-Buffer captures recent experiences essential for immediate policy updates, aligning with findings by Nagabandi et al. [14], which highlight the importance of rapid adaptation in dynamic environments. The Slow-Buffer, in contrast, serves as a memory-efficient repository, structured to retain a broader distribution of experiences, thereby supporting policy robustness across a range of states and actions. Inspired by Al-Shedivat et al. [15], who emphasized clustering for continual learning, the Slow-Buffer organizes experiences through clustering to preserve the diversity of learned representations [10, 16, 17]. To bridge these identified knowledge gaps, our study introduces an innovative framework which uniquely combines adaptive smart clustering with a dual-buffer strategy to enhance learning efficiency while maintaining stability in complex, changing environments.
In this study, we introduce a novel RL framework, SODACER, that uniquely combines a clustering-based ER strategy with a dual-buffer system to overcome the challenges of traditional ER methodologies. The Fast-Buffer prioritizes low-bias, high-variance samples that enhance short-term responsiveness to policy changes, while the Slow-Buffer, organized through a self-organizing clustering mechanism, captures a diverse range of experiences for a comprehensive environmental representation. This dual-buffer design helps address the bias-variance trade-off by utilizing both recent and historical experiences, reducing redundancy, and supporting policy generalization. By selectively consolidating experiences, SODACER minimizes memory requirements while preserving essential variations a critical feature for resource-constrained RL applications.
To optimize computational efficiency within this framework, we employ the Sophia optimizer [23], known for its dynamic step adjustments that improve convergence rates. For systems experiencing frequent or severe disturbances, we advise adopting a robust, dynamic optimization solution like ET-DSIWO [24]. This integration enhances the adaptability of the SODACER framework in high-dimensional RL applications, further contributing to efficient gradient-based optimization. By combining SODACER with the Sophia optimizer and validating it on a real-world public health model, this study contributes a scalable, robust RL approach for complex, dynamic systems, offering new pathways for effective control in high-dimensional, constraint-driven applications.
We validate our methodology by applying SODACER to the optimal control of a compartmental model for Human Papillomavirus (HPV) transmission dynamics. Given the complex, nonlinear nature of HPV transmission, with inherent state and input constraints, this model serves as a relevant test case for adaptive control policies in a constrained environment. We employed the model and system parameters that had been established previously in our earlier work. Our preceding publication has comprehensively examined and detailed this disease model and convergence study [25]. Our approach addresses the unmet need for RL-based public health strategies, aiming to minimize infection rates and intervention costs through safe, data-driven policies. Prior studies by Saldaña et al. [18] and Malik et al. [19, 20] have underscored the challenges of controlling HPV spread, while Brown et al. [21] highlighted the cost-effectiveness of vaccination programs. This work builds on these insights, demonstrating the potential of safe RL methods for optimizing health policy.
The main contributions of this study are as follows.
  • Leveraged a self-organizing adaptive clustering mechanism in SODACER to dynamically remove redundant experiences, improving memory efficiency and accelerating learning convergence.
  • Developed a dual-buffer architecture combining a fast buffer for immediate instances and a slow buffer for extended experience diversity, effectively managing the bias-variance trade-off.
  • Incorporated CBFs to enforce state and input constraints, ensuring safe optimized control policies in complex environments.
  • Integrated Sophia optimizer for rapid convergence, supporting scalable reinforcement learning in nonlinear systems.
  • For effective control of HPV transmission within public health frameworks, the proposed approach, demonstrates capability to manage infection rates under operational constraints and provides strategic insights for health interventions.
  • Proposed a framework that is inherently scalable and generalizable effectively across a broad range of optimal control problems.
To guide the reader, the paper is organized thematically. Following this introduction, we establish the foundational problem definition and system dynamics. Building on this foundation, we formulate the safe optimal control problem under state constraints. The subsequent sections present our methodological core: first, an online reinforcement learning algorithm using neural networks; then, an examination of experience replay as a reinforcement learning enhancement. This leads to the introduction of our novel SODACER approach for experience replay in dynamic settings. Next, we detail the proposed control mechanism for nonlinear systems. The framework is then validated through a case study using an HPV model, with numerical results showcasing the efficacy of combining the Sophia optimizer with SODACER. A comparative analysis of experience replay methods further contextualizes the contribution. The paper concludes by synthesizing key insights and discussing applications.
Problem Definition and System Dynamics
We consider a category of nonlinear continuous-time systems, modeled in affine form and subject to potential uncertainties. The general formulation for such systems is:
$\dot{x}(t)=f(x(t))+g(x(t))u(t),$ (1)
where $x(t)\in\mathbb{R}^{p}$ represents the system state vector, and $u(t)=[u_{1}(t),u_{2}(t),\dots,u_{q}(t)]^{T}\in\mathbb{R}^{q}$ denotes the vector of control inputs, which is subject to defined constraints [22]. The control input is restricted to the set $u(t)\in\mathcal{V}\subseteq\mathbb{R}^{q}$, where the admissible control set is given by $\mathcal{V}=\{u(t)\in\mathbb{R}^{q}\ |\ |u_{i}(t)|\leq\kappa_{i},\ i=1,\dots,q\}$, and $\kappa_{i}$ denotes the saturation limit for each control input $u_{i}(t)$. The primary objective is to design a control strategy that guarantees stability for system (1) across all $u\in\mathcal{V}$.
In this setup, the functions $f(x)\in\mathbb{R}^{p}$ and $g(x)\in\mathbb{R}^{p\times q}$ define smooth nonlinear dynamics, which are bounded by the conditions $\|f(x)\|\leq\alpha_{f}\|x\|$ and $\|g(x)\|_{F}\leq\alpha_{g}$, where $\alpha_{f}$ and $\alpha_{g}$ are constants, and $\|\cdot\|_{F}$ denotes the Frobenius norm.
A desired reference point, denoted $x_{r}$, defines the target state, with the tracking error expressed as the deviation from the reference:
$e(t)=x(t)-x_{r}.$ (2)
The key aim of this work is to devise an optimal control law that not only stabilizes the expanded system but also minimizes a specified objective function. More precisely, we seek to minimize the following objective function [26]:
$\mathcal{J}(e(t),u)=\int_{t}^{\infty}e^{-\nu(\tau-t)}\mathcal{L}(\tau)d\tau,$ (3)
where $Q\in\mathbb{R}^{p\times p}$ is a positive definite matrix, $\nu$ is a discount factor, and $\mathcal{L}(\tau)=e^{\top}(\tau)Qe(\tau)+U(u(\tau))$. The term $U(u)$ represents the cost associated with control efforts and is given by:
$U(u)=2\sum_{i=1}^{q}\int_{0}^{u_{i}}\left(\kappa_{i}\Phi_{i}\tanh^{-1}\left(\frac{\theta_{i}}{\kappa_{i}}\right)\right)^{T}d\theta_{i},$ (4)
where $\Phi_{i}>0$ for all $i=1,\dots,q$.
The Hamiltonian associated with this optimal control problem can be expressed as:
$\mathcal{H}(e,u,x,\mathcal{J}_{e})=e^{\top}(t)Qe(t)+U(u)-\nu\mathcal{J}(e)$
$+\mathcal{J}_{e}^{\top}\left(f(x(t))+g(x(t))u(t)\right),$ (5)
where $\mathcal{J}_{e}\triangleq\frac{\partial\mathcal{J}(e)}{\partial e}$ represents the gradient of the objective function with respect to the error $e(t)$.
By applying the stationarity condition $\frac{\partial\mathcal{H}(e,u,x,\mathcal{J}_{e})}{\partial u}=0$, the optimal control input $u^{*}=[u_{1}^{*},u_{2}^{*},\dots,u_{q}^{*}]^{T}$ is derived as:
$u_{i}^{*}=-\kappa_{i}\tanh\left(\frac{1}{2\kappa_{i}\Phi_{i}}g^{\top}(x)\mathcal{J}_{e}^{*}(x)\right).$ (6)
By substituting the optimal controls (6) in (3) the corresponding optimal cost functional is obtained, leading to the Hamilton-Jacobi-Bellman (HJB) equation for the expanded system:
$e^{\top}(t)Qe(t)+U(u^{*})-\nu\mathcal{J}^{*}(e)$
$+\mathcal{J}_{e}^{*\top}\left(f(x(t))+g(x(t))u^{*}(t)\right)=0.$ (7)
This research proposes a novel approach to solving the Hamilton-Jacobi-Bellman (HJB) equation by employing an online reinforcement learning methodology. The solution is approximated through the implementation of a neural network architecture, leveraging the power of machine learning to address this complex mathematical problem.
Theorem 1 The optimal control input $u^{*}$, derived in (6), guarantees that the system (1) remains uniformly ultimately bounded (UUB). Consequently, the tracking error $e(t)$ is also uniformly ultimately bounded.
proof The proof follows the methodology outlined in [5].
The implementation of the online reinforcement learning algorithm, as outlined in the next section, builds upon the formulation of the optimal control problem introduced earlier, providing a practical method for approximating the solution using neural networks.
Safe Optimal Control Problem
Consider a nonlinear dynamical system characterized in (1), the objective is to determine a control policy $u(t)$ that minimizes a performance criterion , given in equation (4).
In addition to optimizing the cost functional, the control $u(t)$ must satisfy state constraints that ensure the system operates within safe boundaries. To ensure the system adheres to the state constraints in a dynamically responsive manner, we employ the concept of Safe Optimal Control through Control Barrier Functions (CBFs). A Control Barrier Function $h(x)$ provides a measure of “safety” by enforcing constraints on the state trajectory $x(t)$ under the influence of the control input $u(t)$. These constraints are represented as:
$h(x(t))\geq 0,\quad\forall t\in[0,T],$ (8)
where $h:\mathbb{R}^{n}\to\mathbb{R}$ is a smooth function defining the permissible state space. The inequality $h(x(t))\geq 0$ ensures that the state remains within a safe, admissible set throughout the control horizon.
The optimal control problem with state constraints can thus be formally stated as:
Minimize $\mathcal{J}(e(t),u)=\int_{t}^{\infty}e^{-\nu(\tau-t)}\mathcal{L}(\tau)d\tau,$ (9a)
$\text{s.t.}\quad\dot{x}(t)=f(x(t))+g(x(t))u(t),$ (9b)
$\quad\quad\quad\quad\quad h(x(t))\geq 0.$ (9c)
[…]
$\frac{\partial h(x)}{\partial x}f(x(t))+g(x(t))u+\tilde{\alpha}(h(x))\geq 0,\quad\forall x\in\mathbb{R}^{n}.$ (10)
This inequality implies that, as long as $h(x(0))\geq 0$ at the initial time, the system state $x(t)$ will remain in the safe set $\{x\in\mathbb{R}^{n}:h(x)\geq 0\}$ for all $t\geq 0$.
[…]
Minimize $\mathcal{J}(e(t),u)=\int_{t}^{\infty}e^{-\nu(\tau-t)}\mathcal{L}(\tau)d\tau,$ (11a)
$\text{s.t.}\quad\dot{x}(t)=f(x(t))+g(x(t))u(t)),$ (11b)
$\frac{\partial h(x)}{\partial x}(f(x(t))+g(x(t))u)+\tilde{\alpha}(h(x))\geq 0.$ (11c)
The safe optimal control policy $u^{*}(t)$ is derived by solving the above constrained optimization problem, balancing the objectives of minimizing the cost functional $J$ while satisfying the CBF-based safety constraint. By integrating CBFs into the optimization, the resulting policy inherently ensures safety, enabling the system to achieve performance goals without violating the state constraints.
Implementation of the Online Reinforcement Learning Algorithm Using Neural Networks
In order to approximate the objective function $\mathcal{J}(x(t))$ over the compact set $\Omega$, a single-layer adaptive critic neural network (ANN) is employed. This ANN is trained to approximate the value of the objective function as follows:
$\mathcal{J}^{*}(e(t),W^{*})=\varphi^{\top}(e(t))W^{*}+\tilde{\epsilon}(e(t)),$ (12)
where $W(t)\in\mathbb{R}^{l}$ represents the adjustable weight vector of the ANN, $\varphi(e(t))\in\mathbb{R}^{l}$ is the activation function, and $\tilde{\epsilon}(e(t))$ denotes the approximation error of the ANN. The parameter $l$ corresponds to the number of neurons in the hidden layer.
Accordingly, the gradient of $\mathcal{J}^{*}(e(t))$, denoted as $\triangledown_{e}\mathcal{J}$, is expressed as:
$\triangledown_{e}\mathcal{J}^{*}=\triangledown_{e}\varphi^{\top}W+\triangledown_{e}\tilde{\epsilon}$, (13)
where $\triangledown_{e}\varphi\triangleq\dfrac{\partial\varphi(e)}{\partial e(t)}$ and $\triangledown_{e}\tilde{\epsilon}\triangleq\frac{\partial\tilde{\epsilon}(e)}{\partial e(t)}$ represent the derivatives of the activation function and the error, respectively.
Since the error $\tilde{\epsilon}(e(t))$ is not directly measurable, we aim to estimate the weight vector $W$. Hence, the approximations for $\mathcal{J}(e(t))$ and its gradient are rewritten as:
$\hat{\mathcal{J}}(e(t))=\varphi(e(t))^{\top}\hat{W}$, (14)
$\triangledown_{e}\hat{\mathcal{J}}=\triangledown_{e}\varphi^{\top}\hat{W}$, (15)
where $\hat{W}$ denotes the estimated weight vector.
$\mathcal{J^{*}}-\hat{\mathcal{J}}=\varphi^{\top}(t)(W^{*}-\hat{W})+\tilde{\epsilon}(e(t)),$ (16)
therefore,
$\mathcal{J^{*}}=\hat{\mathcal{J}}+\tilde{\delta}+\tilde{\epsilon}(e(t)),$ (17)
[…]
Using the result from equation (6), the control input can then be updated as:
$\hat{u}_{i}=-\kappa_{i}\tanh\left(\frac{1}{2\kappa_{i}\Phi_{i}}g^{\top}(e)\triangledown_{x}\varphi^{\top}\hat{W}\right),\quad i=1,\dots,q.$ (18)
By substituting equations (14), (15), and (18) into the Hamiltonian function in (Problem Definition and System Dynamics), we obtain the following approximation:
[…]
$+$ $\triangledown_{e}\varphi^{\top}\hat{W}\left(f(x(t))+g(x(t))\hat{u}(t)\right).$
[…]
$\min\hskip 2.84544ptE=\frac{1}{2}\hat{\mathcal{H}}^{T}\hat{\mathcal{H}}$, (20)
Here, the objective function presented in (20) is designed to minimize the quadratic Hamiltonian. By minimizing the objective function $E$, we achieve optimized system performance, characterized by reduced control input $\hat{u}$ while ensuring that the weights $\hat{W}$ converge to their optimal values.
Heuristics in Experience replay
The primary contributions introduce a set of advanced heuristics that enhance the effectiveness of Experience Replay for optimal control problems involving a complex dynamical system. These contributions are detailed as follows:
Experience Replay in Reinforcement Learning, a review
Experience Replay (ER) is a cornerstone technique in reinforcement learning, designed to enhance sample efficiency and policy stability by reusing historical interactions. Traditional ER methods, such as uniform sampling [10], face challenges in balancing exploration-exploitation trade-offs and adapting to nonstationary environments. Prioritized Experience Replay (PER) [11] introduced weighted sampling based on temporal-difference errors, but risks overfitting to outliers. Recent advancements integrate clustering and dual-buffer strategies to preserve diversity and relevance in dynamic settings [12, 13, 14]. Effectively deploying Experience Replay in optimal control demands meticulous adjustment of parameters like replay memory size, batch sampling frequency, and the balance between exploration and exploitation. Furthermore, techniques like PER can enhance learning by emphasizing critical transitions that significantly impact learning goals. Building on the foundational concepts and benefits of Experience Replay, we present a series of heuristics designed to refine the memory structure and optimize sampling strategies specifically for complex optimal control scenarios.
Building on the foundational concepts and benefits of Experience Replay,this paper presents a series of heuristics designed to refine the memory structure and optimize sampling strategies specifically for complex optimal control scenarios. These heuristics, comprising a dual-buffer system, dynamic variance adjustments, and prioritized sampling based on cluster size, aim to enhance the adaptability of control policies while maintaining computational efficiency. Each component leverages the principles of reinforcement learning within the context of high-dimensional and nonlinear control tasks, introducing innovative methods to balance bias and variance, control memory usage, and improve learning outcomes. The following sections detail the key contributions of this approach.
Dual-Buffer Architecture for Enhanced Bias-Variance Balance
In this subsection, we present a dual-buffer system designed to optimize the experience replay mechanism within RL for complex control systems. This architecture addresses the critical challenge of managing bias and variance in policy updates, enhancing both sample efficiency and policy adaptability. Dual-buffer structure consists of two distinct buffers: a Fast Buffer and a Slow Buffer, each serving a unique function in maintaining policy relevance and experience diversity.
The Fast Buffer is a small, adaptive component that prioritizes recent experiences aligned with the current policy $\pi_{\text{curr}}$. New system interactions generate samples, denoted as $S_{\text{new}}=(\mathbf{x}_{t},\mathbf{u}_{t})$, which are immediately stored in the Fast Buffer. By focusing on data that reflects recent agent-environment interactions, the Fast Buffer provides low-bias, high-variance samples, allowing rapid adaptation to evolving dynamics and aligning the agent’s behavior with current environmental conditions.
In contrast, the Slow Buffer functions as a long-term repository, retaining a comprehensive range of experiences from both current and past policies, denoted as $\pi_{\text{curr}}$ and $\pi_{\text{past}}$. As samples in the Fast Buffer age, they transition to the Slow Buffer, preserving broader environmental patterns over time. This process enables the model to generalize across varying conditions by incorporating a more extensive history of interactions, which mitigates overfitting to recent data.
To effectively balance short and long-term experience retention, each new sample is initially placed in the Fast Buffer, providing immediate availability for policy updates. As samples age, they move to the Slow Buffer, contributing to a broader learning base. This dynamic integration process maximizes adaptability by allowing the model to leverage recent, high-variance data from the Fast Buffer while drawing from a rich history in the Slow Buffer.
The dual-buffer architecture enhances learning adaptability and sample efficiency, providing a refined solution to experience replay for RL in dynamic control systems.
Adaptive Cluster Formation and Updating Mechanism
Upon introduction to the Cluster Buffer, each sample $S_{\text{old}}$ undergoes an assessment of its affinity to existing clusters through a membership strength calculation, represented as:
[…]
where $\mathbf{c}_{j}$ denotes the centroid, and $\sigma_{j}^{2}$ represents the variance of cluster $\mathcal{C}_{j}$. This metric enables efficient clustering by ensuring each sample’s integration with the most representative cluster, preventing over-segmentation and enhancing the coherence of clustered data points. The formation of clusters is guided by the computed membership strengths. For any sample with a low similarity to all existing clusters (i.e., $\max(\mu_{\mathcal{C}}(S_{\text{old}}))\leq\Gamma_{th}$), a new cluster $\mathcal{C}_{\text{new}}$ is created in the Cluster Buffer. This new cluster initializes with $\mathbf{c}_{\text{new}}=S_{\text{old}}$ and a variance $\sigma_{\text{new}}=\sigma_{0}=0.02$. Conversely, if a sample fits well within an existing cluster $\mathcal{C}_{j}$, the cluster center $\mathbf{c}_{j}$ is updated to incorporate the new data point:
$\mathbf{c}_{j}=\frac{\mathcal{N}_{j}\mathbf{c}_{j}+S_{\text{old}}}{\mathcal{N}_{j}+1},$ (22)
where $\mathcal{N}_{j}$ denotes the updated sample count in the cluster. This adaptive clustering mechanism ensures that significant data patterns are captured efficiently while maintaining the relevance of established clusters.
Adaptive Cluster Formation for Generalization
To maintain the self-organizing of clusters, a suite of variance management techniques is implemented:
Variance Amplification for Cluster Expansion
For clusters absorbing a new sample $S_{\text{old}}$, the cluster variance $\sigma_{j}$ is increased by a factor $(1+\beta)$ to enhance the cluster’s flexibility in accommodating diverse data:
$\sigma_{j}\leftarrow\sigma_{j}\times(1+\beta).$ (23)
Variance Reduction Considered a Forgetting Factor
A variance reduction mechanism is employed across all clusters, scaling each cluster’s variance $\sigma_{k}$ based on its sample count to maintain generalizability. The adjustment is computed as:
$\sigma_{k}\leftarrow\sigma_{k}\times\sigma_{0}\left(\frac{1}{\rho}\right)\left(1-\frac{\mathcal{N}_{k}}{\sum_{i}\mathcal{N}_{i}}\right),$ (24)
where $\rho$ is a scaling factor, and $\sigma_{0}$ the baseline variance.
Omit of Narrow Clusters for Memory Efficiency
Clusters exhibiting a variance below a specified threshold $\sigma_{\text{th}}$ are pruned from the Cluster Buffer. This selective pruning enhances computational efficiency by removing clusters that no longer represent significant data patterns.
Similar Clusters Merging for Structural Optimization
Clusters with a close spatial proximity, as shown in Figure 1, defined by:
$\|\mathbf{c}_{i}-\mathbf{c}_{j}\|<\gamma\max(\sigma_{i},\sigma_{j}),$ (25)
Now if (25) is satisfied, then overlap between dusters $i$ and $j$ is accured and without loosing the generality, let’s assume $\sigma_{i}\geqslant\sigma_{j}$ then the cluster $\mathcal{C}_{j}$ will be pruned and they are merged to form a consolidated cluster $\mathcal{C}_{\text{merged}}$ with its center recalculated as:
$\mathbf{c}_{\text{merged}}=\frac{\mathcal{N}_{i}\mathbf{c}_{i}+\mathcal{N}_{j}\mathbf{c}_{j}}{\mathcal{N}_{i}+\mathcal{N}_{j}}.$ (26)
$\mathcal{N}_{\text{merged}}=\mathcal{N}_{i}+\mathcal{N}_{j},\quad if(\sigma_{i}>\sigma_{j})\longrightarrow\sigma_{\text{merged}}=\sigma_{i},$ (27)
This merging prevents redundancy and ensures efficient memory usage while preserving representational accuracy.
Figure 1: Gaussian membership functions illustrating a 95% overlap. The blue curve represents the first Gaussian function centered at center zero with a standard deviation one , while the red curve represents the second Gaussian function centered at 0.3 with the standard deviation 0.3. The red cluster is redundant and can be absorbed into the blue cluster for improved efficiency.
Remark 1 The value of the parameter $\gamma$, which quantifies the overlap between two clusters, is derived as follows. The overlap is determined by the membership strength $\mu_{\mathcal{C}_{i}}(c_{j})$ between the centroids $c_{i}$ and $c_{j}$ based on (21). To ensure a sufficient degree of overlap, we impose the following condition,
$\text{max}(\mu_{\mathcal{C}_{i}}(c_{i}),\mu_{\mathcal{C}_{j}}(c_{j})))\geq\bar{\gamma}=0.95\%.$
Taking the natural logarithm, this condition becomes:
$\frac{\|c_{i}-c_{j}\|^{2}}{2\sigma_{i}^{2}}\geq\ln(\bar{\gamma})\quad\Rightarrow\quad\frac{\|c_{i}-c_{j}\|^{2}}{\sigma_{i}^{2}}\leq-2\ln(\bar{\gamma}).$
Thus, the cluster $\mathcal{C}_{j}$ is covered completely with the cluster $\mathcal{C}_{i}$ if:
$\|c_{i}-c_{j}\|\leq\sqrt{-2\ln(\bar{\gamma})}\cdot\sigma_{i}.$
The parameter $\gamma$ is therefore defined as:
$\gamma\triangleq\sqrt{-2\ln(\bar{\gamma})}.$
Substituting $\bar{\gamma}=95\%$, we compute:
$\gamma=\sqrt{-2\ln(0.95)}\approx 0.32.$
The proposed method introduces an advanced dual-buffer system combined with intelligent clustering, dynamic variance adjustments, and efficient memory management. This setup enables an experience replay framework that draws high adaptability from recent data within the Fast-Buffer while leveraging broader generalizability from the Cluster Buffer (Slow-Buffer). These innovations address key challenges in managing the bias-variance trade-off, optimizing sample efficiency, and enhancing memory use in RL. By advancing adaptability and memory efficiency, particularly in clustered data storage which requires significantly less memory than storing all experiences this approach pushes the boundaries of experience replay, equipping reinforcement learning algorithms with the capacity to handle complex, high-dimensional control environments more effectively and reliably.
Proposed Mechanism for Optimal Control in Nonlinear Systems
In this section, a novel approach is introduced to optimal control of nonlinear systems by integrating an online RL algorithm into a neural network framework. This innovative architecture leverages the strengths of the Sophia optimizer in conjunction with a meticulously designed experience replay mechanism, thereby enhancing both learning efficiency and system stability. By embedding Lagrange multipliers within the neural network, proposed methodology facilitates their representation as linear combinations of network weights and activation functions. This structure enables dynamic adaptation of control variables based on current system states, thereby improving responsiveness to changes in complex environments. The systematic workflow of our proposed method is visualized in Figure 2.
The detailed steps of proposed proposed algorithm, including gradient derivation, moment estimates, and error minimization, are comprehensively outlined, providing a systematic approach to achieving optimal control in the nonlinear system. The following subsections present the mathematical foundations and mechanisms integral to this approach.
Adaptive Control Mechanism through Reinforcement Learning
The adaptive control mechanism uses gradient calculations to dynamically adjust control variables based on real-time state information. By leveraging experience replay, which stores interactions in a replay buffer, the algorithm can sample past experiences to improve learning efficiency and reduce temporal correlations in updates.
The experience replay system integrates a dual-buffer approach to enhance sample efficiency and policy adaptability. The Fast-Buffer captures recent experiences for immediate responsiveness, while the Slow-Buffer, organized through a clustering technique, provides diverse and comprehensive coverage of the environment.
This approach enables a strategic balance between immediate adaptability and long-term (clustered) learning, supporting the development of effective control policies for complex, high-dimensional nonlinear systems. The framework advances RL in control applications and offers a scalable solution for generating adaptive policies in dynamic environments.
Algorithm 1 Control System Optimization using Sophia+SODACER
1:Define Problem Components and Initialize Hyper Parameters:
2:a) System dynamics: is defined in equation(1);
3:b) objective function: is defined in equation (3);
4:c) Approximated Hamiltonian equation: is defined in Equation (19);
5:d) Error function: is defined in equation (20);
6:e) Initialize state variables: $x=x_{0}$ ;
7:f) Initialize weights: $W=W_{0}$;
8:h) Initialize moment estimates: $m_{0}=0$ and $v_{0}=0$;
9:i) Set hyperparameters: $\beta,\nu,\rho,\beta_{1},\beta_{2},\epsilon_{0},\eta,\Gamma,\sigma_{0},\gamma$;
10:g) Cluster Buffer: $\mathcal{C}=\{\}$, a collection of clusters for experience replay;
11:while stopping criteria not met over time do
12:  Compute Control Law $\mathbf{u}_{t}$ based on $\mathbf{x}_{t}$ and $W_{t}$;
13:  Simulate system dynamics in equation(1);
14:  Generate a new sample $S_{new}=(\mathbf{x}_{t},\mathbf{u}_{t})$ from the latest system experience and add $S_{new}$ into the Fast-Buffer;
15:  Extract the oldest sample from the Fast-Buffer and push it into $S_{old}$;
16:  Compute the membership strength of $S_{old}$ in all existing clusters $\mathcal{C}_{j}$, $\mu_{\mathcal{C}_{j}}(S_{old})$, using (21).
17:  where $\mathbf{c}_{j}$ is the center and $\sigma_{j}^{2}$ is the standard deviation of cluster $\mathcal{C}_{j}$;
18:  if for all clusters $\max(\mu_{\mathcal{C}}(S_{old}))\leq\Gamma_{th}$ then
19:   Add a new cluster $\mathcal{C}_{\text{new}}$ in the Cluster Buffer $\mathcal{C}$ based on $S_{old}$; $\mathbf{c}_{new}=S_{old}$ and $\sigma_{new}=\sigma_{0}=0.02$;
20:  else
21:   Update the cluster $\mathcal{C}_{j}$ with the maximum membership $\mu_{\mathcal{C}}(S_{old})$ by adding $S_{old}$ to $\mathcal{C}_{j}$; $C_{j}=\frac{\mathcal{N}_{j}C_{j}+S_{old}}{\mathcal{N}_{j}+1}$, $\mathcal{N}_{j}=\mathcal{N}_{j}+1$;
22:  end if;
23:  Variances Amplification: For the cluster $\mathcal{C}_{j}$ that covers $S_{old}$, increase variance $\sigma_{j}\leftarrow\sigma_{j}\times(1+\beta)$;
24:  Variances Reduction (Forgetting Factor): For all clusters $\mathcal{C}_{k}$, reduce variance $\sigma_{k}\leftarrow\sigma_{k}\times\sigma_{0}(\frac{1}{\rho})(1-\frac{\mathcal{N}_{k}}{\sum(\mathcal{N}_{i})})$;
25:  Omitting Narrow Clusters:
26:  if for all clusters $\sigma_{\mathcal{C}_{k}}\leq\sigma_{th}$ then
27:   delete the cluster $\mathcal{C}_{k}$;
28:  end if
29:  Similar Cluster Merge:
30:  if for all clusters $\|C_{i}-C_{j}\|<\gamma\times\max(\sigma_{i},\sigma_{j})$ then
31:   merge the clusters $\mathcal{C}_{i}$ and $\mathcal{C}_{j}$ into $\mathcal{C}_{merged}$; $C_{\text{merged}}=\frac{\mathcal{N}_{i}C_{i}+\mathcal{N}_{j}C_{i}}{\mathcal{N}_{i}+\mathcal{N}_{j}}$,  $\mathcal{N}_{merged}=(\mathcal{N}_{i}+\mathcal{N}_{j})$, and $\sigma_{merged}=\max(\sigma_{i},\sigma_{j})$;
32:  end if
33:  Generate mini-batches $\mathcal{B}$: by sampling from all clusters in $\mathcal{C}$, with priority on clusters containing more samples; $\mathcal{B}\leftarrow$ { Fast-Buffer, Center of Clusters in Slow-Buffer, Random Samples of Slow-Buffer Cluster according to the clusters size };
34:  while weight convergence is achieved: $\varphi(t)^{\top}|W_{t+1}-W_{t}|\leq\delta$ do
35:   Compute gradients $\nabla_{W}J(W)$ on the mini-batches:
$\nabla_{W}J(W)=\frac{1}{|\mathcal{B}|}\sum_{i\in\mathcal{B}}\nabla_{W}L(\mathbf{x}_{i},\mathbf{u}_{i})$
36:   Update first moment estimate $m_{t}$ on the mini-batches
$m_{t}=\beta_{1}m_{t-1}+(1-\beta_{1})\nabla_{W}J(W)$
37:   Update second moment estimate $v_{t}$ on the mini-batches
$v_{t}=\beta_{2}v_{t-1}+(1-\beta_{2})\nabla_{W}J(W)$
38:   Correct bias in moment estimates
$\hat{m}_{t}=\frac{m_{t}}{1-\beta_{1}^{t}},\quad\hat{v}_{t}=\frac{v_{t}}{1-\beta_{2}^{t}}$
39:   Update weights $W$ using the Sophia optimizer:
$W_{t+1}=W_{t}-\eta\frac{\hat{m}_{t}}{\sqrt{\hat{v}_{t}}+\epsilon_{0}}$
40:  end while
41:  Update $t\leftarrow t+1$;
42:end while;
43:Output and analyze results based on the updated state variables and control inputs over time;
Problem statement for Human Papillomavirus as a case study
This section provides a brief overview of the transmission dynamics of HPV and the objective function utilized in our previous study [25]. For further details, please consult our earlier work. The model categorizes the population $N(t)$ over time $t$ by gender and infection status, providing a comprehensive framework for analyzing HPV transmission and vaccination impacts. Specifically, the female population $N_{f}(t)$ is segmented into four distinct compartments: unvaccinated susceptible individuals $S_{f}(t)$, vaccinated susceptible individuals $V_{f}(t)$, and infectious females who are either unaware $U_{f}(t)$ or aware $I_{f}(t)$ of their infection status.
The objective is to minimize the spread cost function while adhering to the constraints imposed by disease dynamics, budgetary limitations, and control constraints. Consequently, the optimal control problem is formulated as follows:
$\text{Minimize}\penalty 10000\ J_{\text{spread}}$ $=\int_{0}^{T}A_{0}(I_{m}+I_{f}+U_{f})dt$ (28)
s. t. $\dot{U}_{f}=\left((1-U_{f}-I_{f}-V_{f})+\epsilon V_{f}\right)(1-p)\beta_{m}I_{m}-(\gamma_{f}+\alpha(t)+\mu_{f})U_{f},$
$\dot{I}_{f}=\left((1-U_{f}-I_{f}-V_{f})+\epsilon V_{f}\right)p\beta_{m}I_{m}+\alpha(t)U_{f}-(\gamma_{f}+\mu_{f})I_{f},$
$\dot{V}_{f}=w_{1}(t)\mu_{f}+u_{1}(t)\left(1-U_{f}-I_{f}-V_{f}\right)-\epsilon\beta_{m}V_{f}I_{m}-(\mu_{f}+\theta)V_{f},$
$\dot{I}_{m}=\left(\beta_{f}U_{f}+\widetilde{\beta}_{f}I_{f}\right)\left((1-I_{m}-V_{m})+\epsilon V_{m}\right)-(\gamma_{m}+\mu_{m})I_{m},$
$\dot{V}_{m}=w_{2}(t)\mu_{m}-(\beta_{f}U_{f}+\widetilde{\beta}_{f}I_{f})\epsilon V_{m}+u_{2}(t)\left(1-I_{m}-V_{m}\right)-(\mu_{m}+\theta)V_{m},$
$\dot{J}_{\text{cost}}=\frac{1}{2}\left[A_{1}(w_{1}(t)^{2}+w_{2}(t)^{2})+A_{2}(u_{1}(t)^{2}+u_{2}(t)^{2})+A_{3}\alpha(t)^{2}\right],$
$0\leq U_{f},I_{f},V_{f},I_{m},V_{m}\leq 1,$
$J_{\text{cost}}(0)=0,\quad J_{\text{cost}}(T)=J_{\text{Max}},$
$0\leq w_{1},w_{2}\leq 1,\quad 0\leq u_{1},u_{2}\leq u_{\text{Max}},\quad 0\leq\alpha\leq\alpha_{\text{Max}}.$
The optimal control problem, subject to the specified constraints, ensures that vaccination and screening strategies are effectively implemented within the feasible region. This approach confirms the equilibrium points and assesses the stability of the system.
The activation functions $\varphi(t)$ can be defined as a vector of polynomial functions derived from the system states. For example:
$\varphi(t)=$ $[U_{f},I_{m},V_{m},I_{f},V_{f},U_{f}^{2},I_{f}^{2},V_{f}^{2},I_{m}^{2},V_{m}^{2},$ (29)
$U_{f}\cdot I_{m},U_{f}\cdot V_{m},I_{f}\cdot I_{m},I_{f}\cdot V_{m},V_{f}\cdot I_{m},V_{f}\cdot V_{m},$
$U_{f}^{2}\cdot I_{m},U_{f}^{2}\cdot V_{m},I_{f}^{2}\cdot I_{m},I_{f}^{2}\cdot V_{m},I_{m}^{2}\cdot U_{f},I_{m}^{2}\cdot I_{f},I_{m}^{2}\cdot V_{f},$
$V_{f}^{2}\cdot I_{m},V_{f}^{2}\cdot V_{m},V_{m}^{2}\cdot U_{f},V_{m}^{2}\cdot I_{f},V_{m}^{2}\cdot V_{f}].$
This vector encompasses direct representations of the system states, second-order terms that capture interactions and nonlinearities, as well as cross-terms that account for interactions between females and males. The composition can be tailored to meet specific problem requirements and characteristics.
Table 1: Parameter values for the control model (28)
Parameter Range Mean Value Units Source
$1-\epsilon$ $[0.8,1]$ 0.9 adimensional [18]
$1/\theta$ $[5,15]$ 10 year [18]
$\beta_{m}$ $[0.05,5]$ 4.0 year-1 [18]
$\beta_{f}$ $[0.05,5]$ 4.0 year-1 [18]
$\widetilde{\beta}_{f}$ $[0.025,2.5]$ 2.0 year-1 [25]
$1/\gamma_{f}$ $[0.83,2]$ 1.3 year [18]
$1/\gamma_{m}$ $[0.33,1.2]$ 0.6 year [18]
$p$ $[0,1]$ 0.2 adimensional [25]
$1/\mu_{f}$ $[15,50]$ 30 year [25]
$1/\mu_{m}$ $[15,50]$ 30 year [25]
$A_{1}/A_{0}$ - 0.5 USD/year [25]
$A_{2}/A_{0}$ - 0.2 USD/year [25]
$A_{3}/A_{0}$ - 0.4 USD/year [25]
$u_{\text{max}}$ - 3 year-1 [25]
$\alpha_{\text{max}}$ - 3 year-1 [25]
$J_{\text{max}}$ - 200 USD [25]
Figure 2: Sequential Workflow of the Proposed Approach
Numerical Results of Optimal Control for the HPV Model Using Sophia Optimizer with SODACER
This section presents a thorough assessment of the proposed methodology. Initially, an analysis of the system is conducted in the absence of control mechanisms, utilizing parameter values as specified in Table 1. The resulting simulation outcomes for the system states are depicted in Figure 3 (upper panel). Subsequently, fixed control values are introduced, and the dynamic evolution of system state variables under the influence of our proposed approach starting from predefined initial conditions—is illustrated in Figure 3 (lower panel). To further validate the adaptability of the proposed method, we consider random initial conditions and evaluate its performance over 200 iterations. The results are depicted in Figure 4 and Figure 5, showing the spectrum of state variables and control signals, respectively. In these figures, the blue line represents the average values over the 200 iterations, providing insights into the consistency and effectiveness of the proposed approach.
Figure 3: System states of the HPV model: top panel shows states without control over time, and bottom panel displays states with constant controls ($u_{1}=0.5$, $u_{2}=0.2$, $w_{1}=0.2$, $w_{2}=0.1$, $\alpha=0.2$) over time.
Figure 4: Spectral representation of HPV system states using the proposed approach, based on 200 simulation runs.
Figure 5: Spectral representation of control signals using the proposed approach, based on 200 simulation runs.
Comparative Analysis of Experience replay methods
To evaluate the proposed SODACER-Sophia framework, a comparative analysis was conducted against two baseline methods: Random Experience Replay (RER) and Clustering-Based Experience Replay (CBER). The evaluation considered five distinct cost functions ($f_{1}$$f_{5}$), each representing some control scenario, as summarized in Table 3. These scenarios were designed to analyze the effectiveness of various control strategies for the HPV model by altering the combinations of control variables in the optimization problem (28).
In Scenario 1 ($f_{1}$), vaccination rates for pre-sexually active individuals ($w_{1},w_{2}$), sexually active individuals ($u_{1},u_{2}$), and the female screening rate ($\alpha$) were set to zero. Scenario 2 ($f_{2}$) focused on the vaccination rates of sexually active individuals ($u_{1},u_{2}$) alongside the screening rate ($\alpha$), emphasizing the impact of targeting actively transmitting populations in reducing HPV prevalence. Scenario 3 ($f_{3}$) isolated the vaccination rates for sexually active individuals ($u_{1},u_{2}$), assessing their individual contribution to transmission mitigation. Scenario 4 ($f_{4}$) concentrated solely on the screening rate ($\alpha$), highlighting the critical role of detection and early intervention strategies in controlling HPV spread. Lastly, Scenario 5 ($f_{5}$) utilized all control variables dynamically, demonstrating the effectiveness of a fully integrated control strategy in optimizing HPV intervention outcomes. Table 2 presents a summary of scenarios, including their corresponding objective functions and activated control variables.
The performance of each experience replay method was evaluated across the five scenarios, maintaining consistent hyperparameters such as learning rate, batch size, and buffer capacity to ensure a fair comparison. We employed the Friedman rank test [27], as summarized in Table 3, was employed to systematically compare the effectiveness of the methods in reinforcement learning-based control optimization. RER, serving as the baseline, relies on random sampling of experiences from the replay buffer. While simple and widely adopted, it lacks mechanisms to prioritize critical experiences, resulting in suboptimal learning efficiency and inadequate handling of the bias-variance trade-off. CBER improves upon RER by grouping experiences into clusters based on similarity metrics, which enhances sample relevance and diversity. However, its static clustering approach limits adaptability, particularly in dynamic environments where system states evolve rapidly. By contrast, SODACER employs a dual-buffer architecture with a self-organizing clustering mechanism, dynamically balancing recent experiences (low-bias, high-variance) with aggregated historical experiences (low-variance, high-relevance). This design allows SODACER to address the bias-variance trade-off effectively, promoting efficient learning while maintaining adaptability to system dynamics.
The performance results, summarized in Table 3, demonstrate that SODACER-Sophia consistently outperformed RER and CBER across all metrics. For instance, in Scenario 5, which leverages the full set of controls dynamically, SODACER-Sophia achieved the best convergence rate and highest sample efficiency, highlighting its ability to optimize control strategies effectively. Figure 6 further illustrates the average results across 200 runs, showing SODACER-Sophia’s superior convergence performance compared to RER and CBER. Additionally, Figure 7 provides a detailed comparison of the performance spectrum across 200 runs for the value of the cost function in Equation (28). The results reveal that SODACER-Sophia exhibited the narrowest gap between the worst and best cases, indicating robust and reliable performance. In contrast, RER and CBER showed greater variability, with mean convergence values of approximately 3.5 and 0.2, respectively.
By dynamically managing experiences through its dual-buffer and clustering mechanisms, SODACER-Sophia minimizes redundant samples and accelerates learning, demonstrating clear advantages in efficiency, and adaptability. These results emphasize the potential of the proposed method for optimizing control strategies in complex and dynamic systems like the HPV model.
Table 2: Summary of Scenarios with Corresponding Objective Functions and Activated Control Variables.
Objective Functions $w_{1}$ $w_{2}$ $u_{1}$ $u_{2}$ $\alpha$
$f_{1}$ (Scenario 1) TrueTrue TrueTrue FalseFalse FalseFalse FalseFalse
$f_{2}$ (Scenario 2) FalseFalse FalseFalse TrueTrue TrueTrue TrueTrue
$f_{3}$ (Scenario 3) FalseFalse FalseFalse TrueTrue TrueTrue FalseFalse
$f_{4}$ (Scenario 4) FalseFalse FalseFalse FalseFalse FalseFalse TrueTrue
$f_{5}$ (Scenario 5) TrueTrue TrueTrue TrueTrue TrueTrue TrueTrue
Additionally, to evaluate the effectiveness of control strategies for the HPV model, five distinct cost functions were formulated, each representing a specific control scenario. In “Scenario 1", the cost function incorporates the full set of controls in problem (28), including vaccination rates for pre-sexually active females and males ($w_{1}(t),w_{2}(t)$), vaccination rates for sexually active individuals ($u_{1}(t),u_{2}(t)$), and the screening rate for females ($\alpha(t)$). “Scenario 2" focuses on vaccination of pre-sexually active individuals, utilizing only $w_{1}(t)$ and $w_{2}(t)$ to assess the impact of early immunization on reducing HPV prevalence. “Scenario 3" targets vaccination of sexually active individuals by employing $u_{1}(t)$ and $u_{2}(t)$, highlighting the role of immunizing sexually active populations in controlling transmission. “Scenario 4" isolates the screening rate $\alpha(t)$ as the sole control, emphasizing the effectiveness of detection and early intervention strategies. Lastly, in “Scenario 5", all controls ($w_{1}(t),w_{2}(t),u_{1}(t),u_{2}(t),\alpha(t)$) are set to fixed values, providing a baseline for evaluating the dynamic optimization capabilities of the proposed methodology. These scenarios enable a systematic analysis of individual and combined control strategies to optimize HPV intervention.
Table 3: Obejective Functions Results and Friedman Ranks for RER, CBER, and SODACER Architectures
Obejective Functions RER CBER SODACER
Scenario 1 ($f_{1}$) 2.84 3.20 2.73
Scenario 2 ($f_{2}$) 2.43 2.07 1.69
Scenario 3 ($f_{3}$) 2.67 2.32 1.78
Scenario 4 ($f_{4}$) 3.87 3.37 2.89
Scenario 5 ($f_{5}$) 5.47 2.40 1.00
Friedman Rank 2.80 2.20 1
Figure 6: The mean value of cost function through 200 runs with three methods
Figure 7: The spectrum view of cost function with proposed approach through 200 runs with three methods
Conclusion
This study introduced an advanced RL framework that integrates a dual-buffer experience replay mechanism with self-organizing clustering (SODACER) and CBFs to achieve optimal control in nonlinear, constrained problems. The synergy of SODACER with the Sophia optimizer demonstrated outstanding performance, delivering significant enhancements in adaptability, sample efficiency, and convergence speed compared to traditional experience replay methods. Notably, across 200 experimental runs, SODACER + Sophia exhibited minimal variance between best-case and worst-case outcomes. This consistency is attributed to SODACER’s self-organizing clustering mechanism, which streamlines learning by discarding redundant experiences and retaining only the most pertinent data, thereby accelerating convergence.
The application of this methodology to the Human Papillomavirus (HPV) control model highlighted its capability to optimize intervention strategies effectively. It successfully reduced infection rates and intervention costs while adhering to predefined safety constraints, demonstrating its potential for public health optimization. A comparative evaluation using the Friedman test further validated the superior performance of SODACER + Sophia, showcasing notable computational efficiency and adaptability in high-dimensional environments. These results underscored its advantage over existing experience replay techniques, particularly in balancing memory utilization and learning quality.
This innovative framework represents a significant advancement in safe RL practices, offering a scalable and resource-efficient solution for complex optimal control problems. Future research could extend the SODACER + Sophia framework to various domains requiring adaptive and efficient control strategies, including robotics, healthcare, and large-scale system optimization. Through these contributions, this study establishes a robust foundation for leveraging RL in solving dynamic, real-world challenges.
Acknowledgements (not compulsory)
The authors thanks to anonymous referees and editors.
Author contributions statement
All the authors contributed equally to this work